The firm is an all-equity firm with assets worth $512,000 and 64,000 shares outstanding. It plans to borrow $120,000 and use these funds to repurchase shares. The firm’s marginal corporate tax is 21%, and it plans to keep its outstanding debt equal to $120,000 permanently. If the firm plan to repurchase shares at $9 per share, what is the expected per share value of equity for the leveraged firm? Please show your work.

  1. A) $8 per share

  2. B) $10.45 per share

  3. C) $8.23 per share

  4. D) $8.59 per share

  5. E) Repurchase will not be successful because the firm’s offer is too low.
Answer #1

Correct answer: C) $8.23 per share

All equity firm value (VE) = $512,000

Shares outstanding = 64,000

Tax rate (t) = 21%

Debt Issued (D) = $120,000

Share repurchase price = $9

No. of share repurchased = 120,000/9 = 13,333

According to M&M Proposition 1 with taxes: Value of Leveraged Firm (VL) would be:

V = VE + Dit

Vi = 512,000 + 120,000 * 0.21

Vi = $537,200

Value of Equity in Leveraged Firm(E):

E=VI-D

E = 537,200 – 120,000

E = $417,200

No. of share outstanding after repuchase = 64000-13333 = 50667

Thus,

Expected per share value of equity of leveraged Firm:

417, 200 50, 667

= 8.23415635423451

= $8.23

Hope it will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.